THE MICROWAVE SPECTRUM OF ARGON-TROPOLONE

WEI LIN, WALLACE C. PRINGLE, STEWART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459; \underline{THOMAS A. BLAKE}, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC05-76RLO 1830).

Using a pulsed jet Fourier transform microwave spectrometer we have recorded the rotational spectrum of the argontropolone van der Waals complex between 6 and 17 GHz. Eighty-six transitions were observed (29 a-type and 57 b-type) with J′′ values ranging from 2 to 8 and maximum K′′_{a} and K′′_{c} values of 5 and 8, respectively. The transitions were fit with a Watson A-reduction Hamiltonian in a I^{r} representation using the SPFIT program^{a}. The fit ground state constants for argon-tropolone are A = 1080.4365(3) MHz, B = 883.4943(3) MHz, C = 749.0571(2) MHz, Δ_{J} = 2.591(2) kHz, Δ_{JK} = −3.32(1) kHz, Δ_{K} = 5.232(9) kHz, δ_{J} = 0.944(1) kHz, δ_{K} = −0.028(8) kHz. The rms error for the transition set in the fit is 4 kHz. When included in the Hamiltonian, the tunneling - rotation interaction term, F (J_{a}J_{b} + J_{b}J_{a}), did not improve the quality of the fit: proton tunneling in the complex appears to be quenched. The argon coordinates in the principal axis system of tropolone monomer are a = 0.43 Å, b = 0.23 Å, and c = 3.48 Å.

aH. M. Pickett, R. L. Poynter, E. A. Cohen, M. L. Delitsky, J. C. Pearson and H. S. P. Muller J. Quant. Spectrosc. Radiat. Transfer 60, 883 (1998).